# Error Bound Analysis for the Regularized Loss of Deep Linear Neural Networks

### v1

@@ 4.2 Proofs of the Error Bound
This, together with (163) and (173), implies
$\left|c_{i}-\varphi(\sigma_{i}^{*})\right|=\left|\varphi(\sigma_{i}(\bm{W}_{L}))-\varphi(\sigma_{i}^{*})\right|\leq\frac{\delta_{y}}{3},\ \forall i\in[r_{\sigma}].$ (174)
According to (33), there exists $k_{i}\in[d_{\min}]$ such that $\varphi(\sigma_{i}^{*})=y_{k_{i}}$ for each $i\in[r_{\sigma}]$. For each $i\in[r_{\sigma}]$, we compute
$\left|y_{k_{i}}-y_{\pi^{-1}(i)}\right|=\left|\varphi(\sigma_{i}^{*})-y_{\pi^{-1}(i)}\right|\leq\left|\varphi(\sigma_{i}^{*})-c_{i}\right|+\left|c_{i}-y_{\pi^{-1}(i)}\right|\leq\frac{2\delta_{y}}{3},$
where the last inequality uses (172) and (173). This, together with the definition of $\delta_{y}$ in (24), yields $\varphi(\sigma_{i}^{*})=y_{k_{i}}=y_{\pi^{-1}(i)}$ for each $i\in[r_{\sigma}]$. Using this and (82), we have
For each $i\in[r_{\sigma}+1,d_{\min}]$, we note that $\sigma^{*}_{i}=0$. It is trivial to see that the above equation holds. These, together with the definition of $\mathcal{B}$ in (34) and Lemma 12 in the appendix, yields that $(\bm{\sigma}^{*},\bm{\Pi}^{T})\in\mathcal{B}$. Then, we complete the proof.
$\left|\varphi(\sigma_{i}(\bm{W}_{L}))-\varphi(\sigma_{i}^{*})\right|$ $=\left|c_{i}-y_{\pi^{-1}(i)}\right|\leq\|\bm{c}-\bm{\Pi}^{T}\bm{y}\|=\|\bm{\Pi}\bm{c}-\bm{y}\|$
$=\|\bar{\bm{Y}}-\bm{Y}\|_{F}\leq\frac{2^{L-1}}{\sqrt{\lambda}\sigma_{\min}^{*{L-1}}}\sqrt{\eta_{3}^{2}+\eta_{4}^{2}}\|\nabla G(\bm{W})\|_{F},$ (175)
where the last inequality uses (165). By (76) and the continuity of $\varphi^{\prime}(x)$, Lemma 7 holds, and thus we conclude that there exists a positive constant $\delta_{2}>0$ such that for all $|x-\sigma_{i}^{*}|\leq\delta_{2}$ and all $i\in[r_{\sigma}]$, we have for all $L\geq 3$,
$\left|\varphi^{\prime}(x)-\varphi^{\prime}(\sigma_{i}^{*})\right|\leq\frac{|\varphi^{\prime}(\sigma_{i}^{*})|}{2}.$ (176)
Note that there exists $\delta_{2}>0$ such that (176) holds when $L=2$ (see Remark 2). Using (83a) and Weyl’s inequality, it holds that
$\left|\sigma_{i}(\bm{W}_{L})-\sigma_{i}^{*}\right|\leq\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\delta_{2}.$ (177)
Applying the mean-value theorem to $\varphi(\cdot)$, there exists $x$ between $\sigma_{i}(\bm{W}_{L})$ and $\sigma_{i}^{*}$ such that
$\left|\varphi(\sigma_{i}(\bm{W}_{L}))-\varphi(\sigma_{i}^{*})\right|=\left|\varphi^{\prime}(x)\right|\left|\sigma_{i}(\bm{W}_{L})-\sigma_{i}^{*}\right|\geq\frac{\left|\varphi^{\prime}(\sigma_{i}^{*})\right|}{2}\left|\sigma_{i}(\bm{W}_{L})-\sigma_{i}^{*}\right|,$ (178)
where the inequality follows from $|x-\sigma_{i}^{*}|\leq\left|\sigma_{i}(\bm{W}_{L})-\sigma_{i}^{*}\right|\leq\delta_{2}$ due to (177) and $\left|\varphi^{\prime}(x)\right|\geq|{\varphi^{\prime}(\sigma_{i}^{*})}/{2}|$ due to (176). This, together with (4.2), yields for all $i\in[r_{\sigma}]$,
Using this and (60) in Lemma 3, we obtain for eacl $l\in[L]$,
$\left|\sigma_{i}(\bm{W}_{l})-\sigma_{i}^{*}\right|\leq\left(\frac{2^{L}\sqrt{\eta_{3}^{2}+\eta_{4}^{2}}}{\sqrt{\lambda}\sigma_{\min}^{*{L-1}}\min_{i\in[r_{\sigma}]}\left|\varphi^{\prime}(\sigma_{i}^{*})\right|}+\frac{3\sqrt{2}(L-1)\sigma_{\max}^{*}}{4\lambda\sigma_{\min}^{*}}\right)\|\nabla G(\bm{W})\|_{F}.$
Then, we complete the proof. ∎
@@ 5 Experimental Results
In this section, we conduct experiments under different settings to validate our theoretical results. Specifically, we employ GD to solve Problem (1) using the PyTorch library. Our codes are implemented in Python on a workstation equipped with $24$ GB of RAM and an AMD Ryzen-7 8845H processor with integrated Radeon 780M Graphics operating at $3.80$ GHz. We terminate the algorithm when the squared gradient norm satisfies $\|\nabla f(\bm{W}^{k})\|_{F}^{2}\leq 10^{-6}$ and the function value change satisfies $|f(\bm{W}^{k})-f(\bm{W}^{k-1})|\leq 10^{-7}$ for all $k=1,2,\dots$, where $\bm{W}^{k}$ is the $k$-th iterate.
@@ 5.1 Linear Convergence to Critical Points
In this subsection, we investigate the convergence behavior of GD to different critical points of Problem (2). In our experiments, we set $d_{L}=20$, $d_{0}=10$, and $d_{l}=32$ for each $l=2,\dots,L-1$. Then, we i.i.d. sample each entry of $\bm{Y}\in\mathbb{R}^{d_{L}\times d_{0}}$ from the standard Gaussian distribution, i.e., $y_{ij}\overset{i.i.d.}{\sim}\mathcal{N}(0,1)$. Moreover, we set the regularization parameters $\lambda_{l}=10^{-4}$ for all $l\in[L]$ and the learning rate $4.5\times 10^{-4}$. To ensure convergence to different critical points, we initialize the weight matrices in the neighborhood of two different critical points of Problem (2). Now, we specify how to construct these critical points as follows. We apply the SVD to $\bm{Y}$ and solve the equation (8) to get its roots. Using these results and 1, we can respectively construct an optimal solution, denoted by $\bm{W}_{\rm opt}^{*}$, and a non-global critical point, denoted by $\bm{W}_{\rm crit}^{*}$, of Problem (2). Then, we set $\bm{W}^{0}=\bm{W}^{*}+0.01\bm{\Delta}$, where $\bm{W}^{*}$ is $\bm{W}_{\rm opt}^{*}$ or $\bm{W}_{\rm crit}^{*}$ and each entry of $\bm{\Delta}$ is i.i.d. sampled from the standard Gaussian distribution. For each initialization, we run GD for solving Problem (2) with different depths $L\in\{2,4,6\}$.
@@
Figure 1: Linear convergence of GD to a critical point of Problem (2). In (a) and (b), $\{\bm{W}^{k}\}$ converges to an optimal solution $\bm{W}_{\rm opt}^{*}$; in (c) and (d), $\{\bm{W}^{k}\}$ converges to a non-optimal critical point $\bm{W}_{\rm crit}^{*}$. $F^{\rm end}$ denotes the training loss at the final step when the stop condition is triggered.
We respectively plot the function value gap $F(\bm{W}^{k})-F(\bm{W}^{\rm end})$ and the squared gradient norm $\|\nabla F(\bm{W}^{k})\|_{F}^{2}$ against the iteration number in Figure 1(a) and (b) (resp., Figure 1(c) and (d)) for $\bm{W}_{\rm opt}^{\rm end}$ (resp., $\bm{W}_{\rm crit}^{\rm end}$). We also report the function values of the final iterate, denoted by $\bm{W}^{\rm end}$, in different settings in Table 1. As observed from Figure 1, and Table 1, GD converges to an optimal solution at a linear rate for solving Problem (2) with different network depths, and similarly, it converges to a non-optimal critical point at a linear rate. This aligns with our linear convergence analysis for all critical points of Problem (2) in 7, which leverages the error bound of Problem (2) and supports 2.
$L=2$ $L=4$ $L=6$
$F(\bm{W}_{\rm opt}^{*})$ $9.2900\times 10^{-3}$ $8.1723\times 10^{-3}$ $9.5264\times 10^{-3}$
$F(\bm{W}^{\rm end}_{\rm opt})$ $9.2985\times 10^{-3}$ $8.2020\times 10^{-3}$ $9.5741\times 10^{-3}$
$F(\bm{W}^{*}_{\rm cri})$ $6.8038\times 10^{-1}$ $6.7906\times 10^{-1}$ $6.8022\times 10^{-1}$
$F(\bm{W}^{\rm end}_{\rm cri})$ $9.2980\times 10^{-3}$ $6.7909\times 10^{-1}$ $6.8027\times 10^{-1}$
@@
From Figure 1(c) and (d) and Table 1, we observe that the convergence behavior of GD to a critical point when $L=2$ differs significantly from those when $L=4$ and $L=6$. This difference arises from the benign global loss landscape of Problem (2) when $L=2$ as shown in [50]. That is when $L=2$, each critical point of Problem (2) is either a global minimizer or a strict saddle point. This, together with the result in [28], implies that GD almost surely avoids strict saddle points and converges to an optimal solution. Consequently, even when the starting point is initialized near a non-optimal critical point, GD will escape from it and eventually converge to a global optimum, although it may require more iterations to do so. In contrast, when $L=4$ and $L=6$, we conjecture that the loss landscape of Problem (2) is not benign, which contains non-optimal local minimizers, to which GD may converge. In the future, we will study the global optimization loss landscape of Problem (2) for different $L$.
@@ 5.2 Convergence Behavior in General Settings
General data input.
In this experiment, we consider general data inputs $\bm{X}_{1},\bm{X}_{2},\bm{X}_{3}$ instead of orthogonal inputs $\bm{X}$. Then, we apply GD to optimize Problem (1) for each of these different data matrices.
Linear networks with bias.
Next, we study the regularized loss of deep linear networks with bias terms as follows:
$\min_{\{\bm{W}_{l},\bm{b}_{l}\}}\;\sum_{i=1}^{N}\left\|\bm{W}_{L}(\bm{W}_{L-1}\ldots(\bm{W}_{1}\bm{x}_{i}+\bm{b}_{1})+\bm{b}_{L-1})+\bm{b}_{L}-\bm{y}_{i}\right\|^{2}+\sum_{l=1}^{L}\lambda_{l}\left(\left\|\bm{W}_{l}\right\|_{F}^{2}+\left\|\bm{b}_{l}\right\|^{2}\right),$
@@
Deep nonlinear networks.
Finally, we study the performance of GD for solving the regularized loss of deep nonlinear networks with different activation functions:
$\min_{\{\bm{W}_{l},\bm{b}_{l}\}}\ \sum_{i=1}^{N}\left\|\bm{W}_{L}\sigma\left(\bm{W}_{L-1}\ldots\sigma\left(\bm{W}_{1}\bm{x}_{i}+\bm{b}_{1}\right)+\bm{b}_{L-1}\right)+\bm{b}_{L}-\bm{y}_{i}\right\|_{F}^{2}+\sum_{l=1}^{L}\lambda_{l}\left(\left\|\bm{W}_{l}\right\|_{F}^{2}+\left\|\bm{b}_{l}\right\|_{F}^{2}\right),$
@@
For the above three different settings, we plot the function value gap $F(\bm{W}^{k})-F(\bm{W}^{*})$ against the iteration number in Figure 2. It is observed from Figure 2 that GD converges to a solution at a linear rate across these settings. This consistent behavior leads us to conjecture that the error-bound condition may hold for deep networks in more general scenarios. Additionally, we observe that the number of iterations required to meet the stopping criterion generally increases as the condition number of the input data becomes larger. Exploring this phenomenon presents an interesting direction for future research.
@@ 6 Conclusions
In this paper, we studied the regularized squared loss of deep linear networks (i.e., Problem (2)) and proved its error bound, a critical regularity condition that characterizes local geometry around the critical point set. This result is not only theoretically significant but also lays the foundation for establishing strong convergence guarantees for various methods for solving Problem (2). To establish the error bound, we explicitly characterized the critical point set of (2) and developed new analytic techniques to show the error bound, which may be of independent interest. Our numerical results across different settings provide strong support for our theoretical findings. One future direction is to extend our analysis to Problem (1) with more general data input $\bm{X}$ and loss functions. Another interesting direction is to investigate the regularized loss for deep nonlinear networks.
Supplementary Material
The appendix is organized as follows. In Appendix A, we provide supplementary results and proofs for Section 3. In Appendix B, we show that 2 is a necessary condition for the error bound to hold for Problem (2). In Appendix C, we provide a unified framework to establish linear convergence under the error bound. Finally, we present auxiliary results to prove our main results for completeness.
Appendix A Supplementary Results and Proofs for Section 3
Lemma 5.
Let $a>0$ be a constant, $\bm{\Sigma}\in\mathbb{R}^{n\times n}$ be a diagonal matrix, and $\bm{Q}\in\mathcal{O}^{n}$ be an orthogonal matrix. Then, if $\|\bm{\Sigma}-a\bm{I}\|\leq a/2$, it holds that
$\|\bm{Q}\bm{\Sigma}^{2}-\bm{\Sigma}^{2}\bm{Q}\|_{F}\geq a\|\bm{Q}\bm{\Sigma}-\bm{\Sigma}\bm{Q}\|_{F}.$
@@ Appendix A Supplementary Results and Proofs for Section 3
For ease of exposition, let $\bm{\Delta}:=\bm{\Sigma}-a\bm{I}$. We compute
$\|\bm{Q}\bm{\Sigma}^{2}-\bm{\Sigma}^{2}\bm{Q}\|_{F}$ $=\|\bm{Q}\left(\bm{\Delta}+a\bm{I}\right)^{2}-\left(\bm{\Delta}+a\bm{I}\right)^{2}\bm{Q}\|_{F}=\left\|\bm{Q}\bm{\Delta}^{2}-\bm{\Delta}^{2}\bm{Q}+2a\left(\bm{Q}\bm{\Delta}-\bm{\Delta}\bm{Q}\right)\right\|_{F}$
where the second inequality follows from $\|\bm{Q}\bm{\Delta}^{2}-\bm{\Delta}^{2}\bm{Q}\|_{F}=\|\bm{\Delta}(\bm{\Delta}\bm{Q}-\bm{Q}\bm{\Delta})+(\bm{\Delta}\bm{Q}-\bm{Q}\bm{\Delta})\bm{\Delta}\|_{F}\leq 2\|\bm{\Delta}\|\|\bm{Q}\bm{\Sigma}-\bm{\Sigma}\bm{Q}\|_{F}$ and $\bm{Q}\bm{\Delta}-\bm{\Delta}\bm{Q}=\bm{Q}\bm{\Sigma}-\bm{\Sigma}\bm{Q}$, and the last inequality uses $\|\bm{\Delta}\|\leq a/2$. ∎
Proof of 3.
(i) According to $L=2$, (46), and (50), we have $\mathcal{W}_{\bm{0}}=\{(\bm{0},\bm{0})\}$. Then, we have
$\mathrm{dist}^{2}(\bm{W},\mathcal{W}_{\bm{0}})=\|\bm{W}_{1}\|_{F}^{2}+\|\bm{W}_{2}\|_{F}^{2}.$ (179)
Without loss of generality, assume that $\|\bm{W}_{1}\|_{F}\geq\|\bm{W}_{2}\|_{F}$. Now, we compute
$\frac{1}{2}\nabla_{\bm{W}_{1}}G(\bm{W})=\bm{W}_{2}^{T}\left(\bm{W}_{2}\bm{W}_{1}-\sqrt{\lambda}\bm{Y}\right)+\lambda\bm{W}_{1},$ (180)
$\frac{1}{2}\nabla_{\bm{W}_{2}}G(\bm{W})=\left(\bm{W}_{2}\bm{W}_{1}-\sqrt{\lambda}\bm{Y}\right)\bm{W}_{1}^{T}+\lambda\bm{W}_{2}.$ (181)
Using (180), we have
$\frac{\sqrt{\lambda}}{2}\|\nabla_{\bm{W}_{1}}G(\bm{W})\|_{F}$ $\geq\left\|\lambda\bm{W}_{2}^{T}\bm{Y}-\lambda^{\frac{3}{2}}\bm{W}_{1}\right\|_{F}-\sqrt{\lambda}\|\bm{W}_{2}^{T}\bm{W}_{2}\bm{W}_{1}\|_{F}$
Moreover, multiplying $\bm{Y}^{T}$ on the both sides of (181), together with the triangular inequality, yields
$\frac{1}{2}\left\|\bm{Y}^{T}\nabla_{\bm{W}_{2}}G(\bm{W})\right\|_{F}\geq\left\|\sqrt{\lambda}\bm{Y}^{T}\bm{Y}\bm{W}_{1}^{T}-\lambda\bm{Y}^{T}\bm{W}_{2}\right\|_{F}-\|\bm{Y}^{T}\bm{W}_{2}\bm{W}_{1}\bm{W}_{1}^{T}\|_{F},$
which implies
$\frac{y_{1}}{2}\left\|\nabla_{\bm{W}_{2}}G(\bm{W})\right\|_{F}\geq\left\|\sqrt{\lambda}\bm{Y}^{T}\bm{Y}\bm{W}_{1}^{T}-\lambda\bm{Y}^{T}\bm{W}_{2}\right\|_{F}-y_{1}\|\bm{W}_{2}\|_{F}\|\bm{W}_{1}\|_{F}^{2}.$
Summing up this inequality with (A) and using the triangular inequality, we obtain
$\ \frac{\sqrt{\lambda}}{2}\|\nabla_{\bm{W}_{1}}G(\bm{W})\|_{F}+\frac{y_{1}}{2}\left\|\nabla_{\bm{W}_{2}}G(\bm{W})\right\|_{F}$
$\geq$ $\ \left(\sqrt{\lambda}\min\left\{\min_{i\in[s_{p_{Y}}]}\left|\lambda-y_{i}^{2}\right|,\lambda\right\}\right)\|\bm{W}_{1}\|_{F}-\left(\sqrt{\lambda}+y_{1}\right)\|\bm{W}_{1}\|_{F}^{3}$
where the second inequality follows from (22), $\lambda\neq y_{i}^{2}$ for all $i\in[p_{Y}]$ due to (75), and $\|\bm{A}\bm{B}\|_{F}\geq\sigma_{\min}(\bm{A})\|\bm{B}\|_{F}$ when $\bm{A}$ is non-degenerate, and the last inequality uses (51). This, together with (179), directly implies (52).
(ii) According to (46) and (50), we have $\mathcal{W}_{\bm{0}}=\{(\bm{0},\bm{0},\dots,\bm{0})\}$. Then, we have
$\mathrm{dist}^{2}(\bm{W},\mathcal{W}_{\bm{0}})=\sum_{l=1}^{L}\|\bm{W}_{l}\|_{F}^{2}.$ (183)
Let $k\in[L]$ be such that $k\in\operatorname{\arg\max}_{l\in[L]}\|\bm{W}_{l}\|_{F}$. Then, we compute
[…]
$=\left(\lambda-\|\bm{W}_{k}\|_{F}^{2L-2}-\sqrt{\lambda}y_{1}\|\bm{W}_{k}\|_{F}^{L-2}\right)\|\bm{W}_{k}\|_{F}\geq\frac{\lambda}{3}\|\bm{W}_{k}\|_{F},$
where the last inequality uses (53). This, together with (183), yields
$\mathrm{dist}^{2}(\bm{W},\mathcal{W}_{\bm{0}})\leq L\|\bm{W}_{k}\|_{F}^{2}\leq\frac{9L}{4\lambda^{2}}\left\|\nabla_{\bm{W}_{k}}G(\bm{W})\right\|_{F}^{2},$
which directly implies (54). ∎
Appendix B Showing the Necessity of 2
Before we proceed, we claim that the error bound of $F$ (see Problem (2)) holds if and only if the error bound of $G$ (see Problem (13)) holds. Indeed, it follows from (ii) in Lemma 1 that the “if” direction holds. Now, it remains to show the “only if” direction. Suppose that the error bound of Problem (2) holds (see (11)) for all critical points. Using the proof setup in Lemma 1, it follows (18) that
$\frac{\lambda}{\lambda_{\max}}\,\mathrm{dist}(\bm{Z},\mathcal{W}_{G})\leq\frac{\lambda}{\sqrt{\lambda_{\max}}}\,\mathrm{dist}(\bm{W},\mathcal{W})\leq\frac{\lambda\kappa_{1}}{\sqrt{\lambda_{\max}}}\,\|\nabla F(\bm{W})\|_{F}\leq\kappa_{1}\,\|\nabla G(\bm{Z})\|_{F},$
where the second inequality follows from (11) and the last inequality uses (3.1). Then, we prove the claim. Based on the error-bound equivalence between $F$ and $G$, establishing the necessity of 2 for $F$ is equivalent to doing so for $G$.
B.1 The Case $L=2$
Lemma 6.
Suppose that $L=2$ and (75) does not hold. For any $\bm{\sigma}^{*}\in\cal A$, the error bound for the critical point set $\widehat{\mathcal{W}}_{\bm{\sigma}^{*}}$ of Problem (13) fails to hold, i.e., for any $\kappa,\delta>0$, there exists $\bm{W}$ satisfying $\mathrm{dist}(\bm{W},\widehat{\mathcal{W}}_{\bm{\sigma}^{*}})\leq\delta$ such that $\|\nabla G(\bm{W})\|_{F}\leq\kappa\cdot\mathrm{dist}(\bm{W},\mathcal{W}_{G})$.
Proof.
@@ B.1 The Case $L=2$
Recall that $\lambda=\lambda_{1}\lambda_{2}$ when $L=2$. Since (75) does not hold and , there exists $i\in[r_{Y}]$ such that $y_{i}=\sqrt{\lambda}$. Now, we define $\bm{W}(t):=\left(\bm{W}_{1}(t),\bm{W}_{2}(t)\right)$ as follows:
@@
$\begin{cases}&\bm{W}_{1}(t)=\bm{Q}_{2}\bm{\Sigma}_{1}(t)\mathrm{BlkDiag}\left(\bm{O}_{1},\dots,\bm{O}_{p_{Y}},\bm{O}_{p_{Y}+1}\right),\\ &\bm{W}_{2}(t)=\mathrm{BlkDiag}\left(\bm{O}_{1}^{T},\dots,\bm{O}_{p_{Y}}^{T},\widehat{\bm{O}}_{p_{Y}+1}^{T}\right)\bm{\Sigma}_{2}(t)\bm{Q}_{2}^{T},\\ &\bm{\Sigma}_{l}(t)=\mathrm{BlkDiag}\left(\mathrm{diag}(\bm{\sigma}^{*}+t\bm{e}_{i}),\bm{0}\right)\in\mathbb{R}^{d_{l}\times d_{l-1}},\forall l=1,2,\\ &\bm{O}_{i}\in\mathcal{O}^{h_{i}},\forall i\in[p_{Y}],\bm{O}_{p_{Y}+1}\in\mathcal{O}^{d_{0}-r_{Y}},\widehat{\bm{O}}_{p_{Y}+1}\in\mathcal{O}^{d_{2}-r_{Y}},\end{cases}$

### v3

@@ 3.2.2 Proof of the Error Bound
Since $\bm{W}^{*}\in\mathcal{W}_{\bm{\sigma}^{*}}$, we have $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\|\bm{W}-\bm{W}^{*}\|_{F}\leq\|\bm{W}-\hat{\bm{W}}\|_{F}+\|\hat{\bm{W}}-\bm{W}^{*}\|_{F}.$ This, together with (119) and (124), implies (113). ∎
Using 3 and the connection of Problems (1) and (12), we are now ready to prove 2.
Proof.
Proof of 2. According to 3, we conclude that for any $\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}\setminus\{\bm{0}\}$, (113) holds for $\mathcal{W}_{\bm{\sigma}^{*}}$, where the constants $\epsilon_{\bm{\sigma}^{*}}$ and $\kappa_{\bm{\sigma}^{*}}$ (see (114) and (115)) depend on $\bm{\sigma}^{*}$. Recalling the definition of $\mathcal{A}$ in (26), and noting that a polynomial has only finitely many roots, it follows that $\mathcal{A}_{\rm sort}$ defined in (29) is finite. Consequently, by (29), we can derive a corresponding result for $\mathcal{W}_{G}$ that does not depend on $\bm{\sigma}^{*}$. In addition, we take the analysis about $\bm{\sigma}^{*}=\bm{0}$ in 2 into consideration. Specifically, by letting
$\kappa:=\max_{\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}}\kappa_{\bm{\sigma}^{*}}\quad\text{and}\quad\epsilon:=\min_{\bm{\sigma}*\in\mathcal{A}_{\rm sort}}\epsilon_{\bm{\sigma}^{*}},$ (125)
we have $\mathrm{dist}(\bm{W},\mathcal{W}_{G})\leq\kappa\|\nabla G(\bm{W})\|_{F}$ for all $\mathrm{dist}(\bm{W},\mathcal{W}_{G})\leq\epsilon.$ Using this and Lemma 1(ii), (9) holds with
$\kappa_{1}={\kappa\lambda}/{\lambda_{\min}}\quad\text{and}\quad\epsilon_{1}={\epsilon}/{\sqrt{\lambda_{\max}}}.$ (126)
@@ 4 Experimental Results
In this section, we conduct experiments under different settings to validate our theoretical results. Specifically, we employ GD to train deep networks using the PyTorch library. We terminate the algorithm when the squared gradient norm satisfies $\|\nabla F(\bm{W}^{k})\|_{F}^{2}\leq 10^{-6}$ and the function value change satisfies $|F(\bm{W}^{k})-F(\bm{W}^{k-1})|\leq 10^{-7}$ for all $k=1,2,\dots$, where $\bm{W}^{k}$ is the $k$-th iterate.
@@ 4.1 Linear Convergence to Critical Points
In this subsection, we investigate the convergence behavior of GD to different critical points of Problem (1). In our experiments, we set $d_{L}=20$, $d_{0}=10$, and $d_{l}=32$ for each $l=2,\dots,L-1$. Then, we i.i.d. sample each entry of $\bm{Y}\in\mathbb{R}^{d_{L}\times d_{0}}$ from the standard Gaussian distribution, i.e., $y_{ij}\overset{i.i.d.}{\sim}\mathcal{N}(0,1)$. Moreover, we set the regularization parameters $\lambda_{l}=10^{-4}$ for all $l\in[L]$ and the learning rate $4.5\times 10^{-4}$. To ensure convergence to different critical points, we initialize the weight matrices in the neighborhood of two different critical points of Problem (1). Specifically, we apply the SVD to $\bm{Y}$ and solve the equation (6) to get its roots. Using these results and 1, we can respectively construct an optimal solution, denoted by $\bm{W}_{\rm opt}^{*}$, and a non-optimal critical point, denoted by $\bm{W}_{\rm crit}^{*}$, of Problem (1). Then, we set $\bm{W}^{0}=\bm{W}^{*}+0.01\bm{\Delta}$, where $\bm{W}^{*}$ is $\bm{W}_{\rm opt}^{*}$ or $\bm{W}_{\rm crit}^{*}$ and each entry of $\bm{\Delta}$ is i.i.d. sampled from the standard Gaussian distribution. For each initialization, we run GD for solving Problem (1) with different depths $L\in\{2,4,6\}$.
@@
Figure 1: Linear convergence of GD to a critical point. In (a) and (b), $\{\bm{W}^{k}\}$ converges to an optimal solution when initialized near an optimal solution; in (c) and (d), $\{\bm{W}^{k}\}$ converges to a non-optimal critical point for $L=4,6$, and to an optimal solution for $L=2$ when initialized near a non-optimal critical point. Here $F^{\rm end}$ denotes the training loss when the stopping criterion is met.
Initialization Objective values $L=2$ $L=4$ $L=6$
Initialization near global optimal solution $F(\bm{W}_{\rm opt}^{*})$ $9.2900\times 10^{-3}$ $8.1723\times 10^{-3}$ $9.5264\times 10^{-3}$
$F(\bm{W}^{\rm end})$ $9.2985\times 10^{-3}$ $8.2020\times 10^{-3}$ $9.5741\times 10^{-3}$
Initialization near non-optimal critical point $F(\bm{W}^{*}_{\rm cri})$ $6.8038\times 10^{-1}$ $6.7906\times 10^{-1}$ $6.8022\times 10^{-1}$
$F(\bm{W}^{\rm end})$ $9.2980\times 10^{-3}$ $6.7909\times 10^{-1}$ $6.8027\times 10^{-1}$
@@
We respectively plot the function value gap $F(\bm{W}^{k})-F(\bm{W}^{\rm end})$ and the squared gradient norm $\|\nabla F(\bm{W}^{k})\|_{F}^{2}$ against the iteration number in Figure 1(a) and (b) (resp., Figure 1(c) and (d)), where $\bm{W}^{\rm end}$ denotes the last iterate of the GD. We also report the function values of the final iterate, in different settings in Table 2. As observed from Figure 1 and Table 2, GD converges to an optimal solution at a linear rate for solving Problem (1) with different network depths, and similarly, it converges to a non-optimal critical point at a linear rate for $L=4,6$. We should point out that Problem (1) only has global optima and strict saddle points when $L=2$, as shown in [6]. Therefore, even if initialized in the vicinity of a non-optimal critical point, i.e., strict saddle point, the algorithm will ultimately converge to a global optimal point almost surely [27]. This aligns with our linear convergence analysis for all critical points of Problem (1) in 6 in Appendix C, which leverages the error bound of Problem (1), and thereby supports 2.
@@ 4.2 Convergence Behavior in General Settings
General data input. We first consider general data inputs $\bm{X}_{1},\bm{X}_{2},\bm{X}_{3}$ instead of orthogonal inputs $\bm{X}$. Then, we apply GD to optimize Problem (1) with a linear network $\bm{W}_{L}\ldots\bm{W}_{1}\bm{X}$ for each of these different data matrices.
Linear networks with bias. We next study the regularized loss of deep linear networks with bias terms:
$\min\limits_{\{\bm{W}_{l},\bm{b}_{l}\}}\;\sum_{i=1}^{N}\left\|\bm{W}_{L}(\bm{W}_{L-1}\cdots(\bm{W}_{1}\bm{x}_{i}+\bm{b}_{1})+\bm{b}_{L-1})+\bm{b}_{L}-\bm{y}_{i}\right\|^{2}+\sum_{l=1}^{L}\lambda_{l}\left(\left\|\bm{W}_{l}\right\|_{F}^{2}+\left\|\bm{b}_{l}\right\|^{2}\right),$
@@
Deep nonlinear networks. Finally, we study the performance of GD for solving the regularized loss of deep nonlinear networks with different activation functions:
$\min\limits_{\{\bm{W}_{l},\bm{b}_{l}\}}\ \sum_{i=1}^{N}\left\|\bm{W}_{L}\sigma\left(\bm{W}_{L-1}\cdots\sigma\left(\bm{W}_{1}\bm{x}_{i}+\bm{b}_{1}\right)+\bm{b}_{L-1}\right)+\bm{b}_{L}-\bm{y}_{i}\right\|_{F}^{2}+\sum_{l=1}^{L}\lambda_{l}\left(\left\|\bm{W}_{l}\right\|_{F}^{2}+\left\|\bm{b}_{l}\right\|_{F}^{2}\right),$
@@
For the above three different settings, we plot the function value gap $F(\bm{W}^{k})-F(\bm{W}^{\text{end}})$ against the iteration number in Figure 2. It is observed from Figure 2 that GD converges to a solution at a linear rate across these settings. This consistent behavior leads us to conjecture that the error-bound condition may hold for deep networks in more general scenarios. Additionally, we observe that the number of iterations required to meet the stopping criterion generally increases as the condition number of the input data becomes larger. Exploring this phenomenon presents an interesting direction for future research.
@@ 5 Conclusions
In this paper, we studied the regularized squared loss of deep linear networks and proved its error bound, a regularity condition that characterizes local geometry around the critical point set. This result is not only theoretically significant but also lays down the foundation for establishing strong convergence guarantees for various methods for solving Problem (1). To establish the error bound, we explicitly characterized the critical point set of (1) and developed new analytic techniques to show the error bound, which may be of independent interest. Our numerical results across different settings provide strong support for our theoretical findings. One future direction is to extend our analysis to deep linear networks with more general data input $\bm{X}$ and loss functions. Another interesting direction is to investigate the regularized loss of deep nonlinear networks.
Appendix A Showing the Necessity of Assumption 2
We claim that the error bound of $F$ (see Problem (1)) holds if and only if the error bound of $G$ (see Problem (12)) holds. Indeed, it follows from (iii) in Lemma 1 that the “if” direction holds. Now, it remains to show the “only if” direction. Suppose that the error bound of Problem (1) holds (see (9)). Using the proof setup in Lemma 1, we have
[…]
Therefore, we prove the claim. Based on the error-bound equivalence between $F$ and $G$, establishing the necessity of Assumption 2 for $F$ is equivalent to doing so for $G$.
A.1 The Case $L=2$
Lemma A.1.
Suppose that $L=2$ and (7) does not hold. For any $\bm{\sigma}^{*}\in\cal A$, the error bound for the critical point set $\widehat{\mathcal{W}}_{\bm{\sigma}^{*}}$ of Problem (12) fails to hold.
@@ A.1 The Case $L=2$
Since (7) does not hold, there exists $i\in[r_{Y}]$ such that $y_{i}=\sqrt{\lambda}$. Now, we define $\bm{W}(t):=\left(\bm{W}_{1}(t),\bm{W}_{2}(t)\right)$ as follows:
@@
$\textstyle\begin{cases}&\bm{W}_{1}(t)=\bm{Q}_{2}\bm{\Sigma}_{1}(t)\mathrm{BlkD}\left(\bm{O}_{1},\dots,\bm{O}_{p_{Y}},\bm{O}_{p_{Y}+1}\right),\\ &\bm{W}_{2}(t)=\mathrm{BlkD}\left(\bm{O}_{1}^{T},\dots,\bm{O}_{p_{Y}}^{T},\widehat{\bm{O}}_{p_{Y}+1}^{T}\right)\bm{\Sigma}_{2}(t)\bm{Q}_{2}^{T},\\ &\bm{\Sigma}_{l}(t)=\mathrm{BlkD}\left(\mathrm{diag}(\bm{\sigma}^{*}+t\bm{e}_{i}),\bm{0}\right)\in\mathbb{R}^{d_{l}\times d_{l-1}},\forall l=1,2,\\ &\bm{O}_{i}\in\mathcal{O}^{h_{i}},\forall i\in[p_{Y}],\bm{O}_{p_{Y}+1}\in\mathcal{O}^{d_{0}-r_{Y}},\widehat{\bm{O}}_{p_{Y}+1}\in\mathcal{O}^{d_{2}-r_{Y}}.\end{cases}$